# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-D Summary
With eye-tracking technology becoming pervasive, as provided in the preceding three sections, three main directions exist in eye-tracked VR. The computer vision and machine learning community addresses eye region segmentation, gaze estimation, and eye event classification tasks to facilitate accurate attention tracking in VR. Deep learning-based end-to-end approaches work successfully when considering different algorithms and methods, particularly for eye region segmentation and gaze estimation. To facilitate eye-tracking technology to be used in an everyday manner in VR, most recently, various researchers and device vendors embedded eye trackers in VR HMDs and provided the community with publicly available datasets to facilitate benchmarking and open science.
Once the eye regions and gaze directions are available, gaze-based interaction, including target selection and text entry through eye movements, predictive gaze analyses, covering gaze and scanpath prediction, and rendering methodologies by keeping the position of the eyes in mind, are the most prominent research directions. In most of these works, real-time working capability and the quality of the eye movement signals are essential to achieve a good interaction experience. Therefore, if the gaze estimation and eye-tracker calibration work reasonably, it is likely that methods relying on the gaze data provide a good usability and user experience.
While real-time working capability and eye-tracking data quality are essential for gaze-based interaction in VR, eye movements have also been researched offline to understand human cognition, visual attention, and perception. As one can assess these human aspects in various domains in VR, the studies have a broader focus. Prominent domains are education and medicine, as there is a pressing shift for online teaching and learning, especially after the COVID-19 pandemic, and it is possible to control environmental and stimulus factors very precisely in VR, which is essential for the medical domain. Regardless of the studied domains, fixation, saccade, blink, pupil diameter, and scanpath statistics are commonly used to assess human cognition, attention, and perception in VR.
@@ V Security and Privacy in Eye Tracking and Implications for Virtual Reality
It is possible to extract and infer useful information about users utilizing eye trackers in VR; therefore, eye tracking is considered a powerful modality in the studies above. We mostly considered works that utilize or are directly applicable to VR setups. However, virtual environments are relatively new, and researchers have not evaluated possibilities with eye movements extensively in VR compared to real-world settings. Despite this, it is known that human eye movements depend on the context and stimulus, and we first argue that as long as users encounter similar stimuli in virtual environments, comparable inferences will be possible, like in the real world. Liebling and Preibusch [4] stated previously that apart from person identification, eye movements can reveal many user attributes such as gender, sexual preferences, body mass index, health status, or tasks, mostly in real-world settings. These would presumably apply to virtual settings, as some of them, such as gender and task predictions, can already be carried out accurately in VR [25, 312, 266]. We foresee many other inference possibilities when such virtual environments and setups are more prevalent in everyday life.
To this end, Xu et al. [313] stated that while the existence of eye-tracking data is vital for utility tasks such as improving the efficiency of VR rendering, it also gives opportunities, especially for companies, for targeted advertising based on such unique characteristics and demographics of users. Considering these, as identification of users might help personalize their VR setups, we first discuss authentication possibilities in Section V-A. However, if the purpose of the task is not authenticating the users, carrying out the same or similar tasks may cause a privacy breach. The National Institute of Standards and Technology (NIST) defines the privacy breach as “the loss of control, compromise, unauthorized disclosure, unauthorized acquisition, or any similar occurrence where a person other than an authorized user accesses or potentially accesses data or an authorized user accesses data for an other than authorized purpose” [314]. According to the NIST’s definition, access and action by authorized entities or data processing with consent does not constitute a privacy breach. However, with increased digitalization, it is possible that adversarial actions and entities may cause privacy breaches. Considering this, previous research studied how to preserve data and user privacy. Section V-B covers these and discusses the methods to preserve privacy when eye-tracking data and systems are the focus in VR.
@@ V-A Eye-based Authentication for VR
Visual interaction and user assistance are two of the most prominent uses of behavioral data, including eye movements, especially in VR applications that work in real time. Apart from those, such behavioral data have also been utilized for authentication purposes. One major challenge for behavioral authentication is that the accuracy of the methods is not very high, which can cause an unpleasant user experience. This section provides an overview of the research on eye and gaze-based authentication in VR.
@@
Pfeuffer et al. [315] studied authentication in VR by utilizing body motion data, including head, hand, and eye movement data for different tasks such as pointing, grabbing, walking, and typing. The authors found that while the best-achieved identification accuracies are near $60\%$, the performance drops when the user group size is increased. The authors also showed that body motions can help with adaptivity in VR. However, they did not use any gaze features apart from the gaze ray. Liebers et al. [316] studied user authentication based on VR spatial movements and reached accuracy up to $90\%$. In addition, in another study, Liebers et al. [317] showed that gazing behaviors and head orientation can authenticate users in VR with almost perfect accuracy. However, the limitation of those studies is the small sample size. Therefore, it is an open question of how the utilized features perform when more users are available and how usable the proposed authentication methods would be. That is why more distinguishing and fine-grained features are needed when authentication through eye movements is desired. One could achieve this using eye-tracking data from high-sampling frequency eye trackers.
@@
When we consider eye-based authentication, there are two prominent directions: iris-based and eye movement-based authentication. On the one hand, while iris textures are like visual fingerprints and help authenticate users with very high accuracies [318, 319], and iris-based authentication can also be used in real-world scenarios such as in airports and border crossings [320]. Especially due to privacy reasons, many commercial HMDs that integrate eye trackers do not provide raw eye images to the users or applications (e.g., HTC Vive Pro Eye [1]). This precaution can somewhat mitigate the problem; however, in the case of iris-textures utilized in VR, data protection issues should be handled carefully, possibly by using encryption schemes [319]. On the other hand, even if the iris textures are not captured or are not accessible with these devices, eye movements will mostly still be provided. Hence, it is plausible to carry out authentication using eye movements in the background over a certain period, depending on the sampling frequencies and granularity of eye movement features, to perform this task. While we do not conduct an extensive review for eye-based authentication, we provide essential studies related to implicit biometric authentication, especially utilizing eye movements in VR, and discuss why they are related to privacy aspects of such data. For more detailed information on eye-based authentication and authentication with different modalities, including eye tracking, we refer the reader to the survey papers by Katsini et al. [26] and Stephenson et al. [321], respectively.
@@
While not necessarily in VR, biometric identification was carried out successfully using oculomotor plant models and eye movements [322]. Task-independent authentication is also possible using eye movements [323]. More for VR, Eberz et al. [324] argued that one could use specific eye movement features for biometric authentication in cheap consumer-level devices during everyday tasks such as reading, writing, and web browsing. While the authors did not use a VR setup, their sampling rate was 50 Hz, even lower than the sampling rates of the eye trackers of today’s consumer-grade HMDs. In follow-up work, Eberz et al. [325] presented a continuous authentication system based on eye movement biometrics and stated that for eye movement-based authentication, a precise calibration should be carried out and effects of light sensitivity and task dependence of eye movements should be considered when designing the authentication systems. In another work, Zhang et al. [326] proposed a continuous authentication scheme based on eye movements for VR headsets and showed that they could continuously authenticate the wearer of the HMD in the background, demonstrating the potential for HMD personalization. Similarly, Zhu et al. [327] used blinking patterns and pupil sizes for user authentication in VR headsets.
@@
While the aforementioned works show the plausibility of biometric authentication based on eye movements, most of them suffer from issues in terms of usability or privacy in the context of VR, which is similar to behavioral authentication using spatial movements. Concerning usability, authentication models based on eye and gaze movements do not have as high accuracies as iris authentication, which could irritate HMD users if errors occur constantly. Furthermore, such models mostly depend on the temporal movement of the gaze regardless of explicit or implicit authentication, which leads to longer authentication times compared to using single iris images. Lohr and Komogortsev [328] stressed the similar implications and partly confirmed that biometric authentication performance using eye movements is comparable to 4-digit PIN entry when eye-tracking signal quality is extremely high, along with very high sampling frequencies (i.e., 1kHz). The authors stated that 5 seconds of eye movement data are needed to achieve such performance. Niitsu and Nakayama [329] also measured measurement time and presentation size of stimuli on biometric authentication, and they found that using a 3-second timeframe maximized the biometric authentication success rates. While the authors did not use a VR setup for this, one can transfer the results to VR as well, as long as the quality of the eye-tracking data and presentation sizes are comparable. Despite these attempts at biometric authentication based on eye movements, Friedman et al. [330] recently showed it is not possible to identify users very accurately based on their eye movements if the considered number of identities is comparably large, which is in line with the research using behavioral data for authentication [315, 316].
@@
These findings imply that while implicit eye-based authentication might not work similarly to PIN-based authentication regarding usability and security, eye movements include patterns that help identify individuals to a certain extent. These data may thus be more useful for personalization, where each user’s data samples are accumulated over time while using VR HMDs and utilized for authentication-related settings such as Two-Factor Authentication (2FA). Since such an authentication process should run continuously in the background, longer waiting times for authentication would not irritate the users, as users would not be actively aware of the ongoing process. This way of authentication will also potentially make the overall VR experience more personalized, as each user’s unique viewing behaviors would be captured.
@@
Several works could support this argumentation. For instance, Lohr et al. [331] proposed a real-time capable architecture for eye movement-based authentication in VR and argued that eye-tracking-based biometrics would become a standard way of authentication for VR. Luo et al. [332] explored the human visual system as a novel authentication method for VR HMDs and proposed OcuLock, which is resistant against impersonation and statistical attacks while maintaining a stable performance over two months, with an increasing error trend when the timespan is increased. Increased error rates in longer timespans pose another prominent challenge than the authentication accuracy itself, especially when the everyday use of such authentication schemes is considered, which requires further research. Furthermore, Friedman and Komogortsev [333] assessed the effectiveness of biometric feature normalization techniques, including real-world eye movement features and synthetic ones. The authors found that the effectiveness of biometric normalization techniques on real-world data depends on the intercorrelation of the features. Mathis et al. [334] used VR to evaluate the usability and security of real-world authentication systems, showing the potential of VR as a testbed for authentication purposes.
@@
Despite several disadvantages, such as the need for consistent tasks, high-sampling frequency eye movement data, and a decrease in authentication performance when more individuals are involved, previous work consistently showed that eye movement patterns represent personal identities. This information is useful if the goal is to authenticate the user in VR or personalize the interaction experience. However, if users do not want to take advantage of these possibilities and yet want to use gaze-based interactions and obtain assistive support, the linkage of eye movements to user identity and characteristics implies privacy risks that should be handled methodologically, meaning that personal patterns associated with eye movements should be either hidden in the data or handled in a privacy-preserving manner. At the same time, one should keep the interaction utility and user experience high during the virtual experience.
@@ V-B Privacy-preserving Eye Tracking for VR
Research in eye-based authentication [26], privacy considerations for eye tracking [4, 23], and privacy risks of data collection in XR [335, pp. 12-14] demonstrate that when authentication is not preferred, it is advisable to preserve the privacy of the individuals and specific attributes of the individuals in the data, in addition to the good data hygiene practices. Recently, what information pupil dilations and gaze information can reveal in XR was also highlighted by Future of Privacy Forum [336], by mentioning concerns on users’ sexual orientation, gender, race, and health. Furthermore, Garrido et al. [337] highlighted data privacy issues in VR, considering threat and defense models and included eye-tracking sensors and data as discussion points. This section covers the methodological research on privacy protection of eye-tracking data.
To provide users with privacy-preserving eye-tracking solutions, researchers have explored a few different approaches applied at various stages of the eye-tracking pipeline. Privacy-preserving methods have been applied to sensor data in the form of eye images, aggregated data in the form of saliency maps and feature vectors, and sample-level gaze position data. Each method transforms or modifies data in various ways to create a trade-off between data privacy and utility that varies by method and intended application of the data.
Recently, formal methods such as differential privacy [338] have attracted the attention of VR and eye-tracking communities. Differential privacy is an established approach, applied across multiple domains, most prominently in the context of databases and aggregate survey data. Differential privacy aims to hide the information on whether an individual’s records exist in a particular database. This privacy mechanism is achieved by adding randomly generated noise according to an $\epsilon$ parameter to query outputs such that including a specific participant in the database does not significantly change the queried function outputs. In $\epsilon$-differential privacy, the smaller the $\epsilon$ is, the more private the mechanism becomes. Since several works focus on differential privacy in the context of eye tracking, we provide the formal definition of the $\epsilon$-differential privacy in the following.
@@
$\epsilon$-Differential Privacy ($\epsilon$-DP) [338, 339]. A mechanism $M$ is considered $\epsilon$-differentially private for all databases $D_{1}$ and $D_{2}$ that differ at most in one element for all $S\subseteq Range(M)$ with;
@@
The eye-tracking community initially utilized standard differential privacy mechanisms, namely Gaussian and Exponential mechanisms, on saliency heatmaps [340] and aggregated eye movement features that were collected from VR reading tasks [25]. These initial works showed the potential of formal methods for the VR domain, even with $\epsilon$ values in the vicinity of $\epsilon=1$; however, they suffer from the correlations in the data that could jeopardize privacy, especially from temporal correlations, as independent noise sampling from the standard mechanisms allows adversaries to reconstruct signals that are very close to the original ones. Despite the correlation issue, both works showed that it is possible to maintain high accuracy in utility tasks while preserving privacy.
@@
Bozkir et al. [312] further iterated on these works and addressed the temporal correlation issue by translating the data into difference signals and using the frequency domain to add the randomly generated noise. The authors noted that apart from the correlation challenge, the privacy-utility trade-off becomes very important as eye-tracking data already consists of a certain amount of noise due to the limitations of sensors and image processing approaches. However, they also indicated that even with these challenges, finding an optimal trade-off is computationally possible in the range of $10<\epsilon<$ 1. Hu et al. [341] took a local differential privacy approach to address the privacy-utility trade-off with a gaze trajectory generation task and argued that their approach could effectively protect privacy for individual users without significant utility decrease. Yet, the privacy noise needed to establish a strong differential privacy guarantee is known to hurt utility, and might end up masking valuable insights from the processed data [342]. To address this, the authors of the previous works that studied differential privacy aimed to calibrate the noise in a way that while privacy is achieved in the reasonable ranges (e.g., $\epsilon<1$), they tried to keep the utility tasks, such as task identification, perform with accuracies double the amount of chance [312]. Yet, preserving privacy in such a way is not trivial.
@@
Based on these prior results, researchers have also pursued alternative privacy guarantees to differential privacy to better understand the range of privacy-utility trade-offs for eye-tracking data. Specifically, David-John et al. explored $k$-anonymity and $k,\gamma$-plausible deniability [343, 344]. These alternative guarantees target the privacy risk of re-identification, and their mathematical formulation is related to the ability to match released data to the original identity in a dataset. Alternative guarantees allow dataset owners to preserve privacy by reducing the risk of linking data to identities and thus mitigating harms related to the inferences that can be made from eye tracking while retaining data utility across different applications. First, $k$-anonymity is a definition of privacy in the context of re-identification from a dataset, and it was proposed by Samarati and Sweeney [345] as follows.
@@
$k$-anonymity. Given a person-specific dataset $D_{1}$, a de-identified dataset $D_{2}$ is $k$-anonymized by privacy process $\mathcal{P}:D_{1}\mapsto D_{2}$ if all released features $\Gamma_{d}=\mathcal{P}(\Gamma)\in D_{2}$ cannot be recognized as $\Gamma$ with probability $>\frac{1}{k}$.
The privacy guarantee is interpreted with a lone privacy parameter $k$ linked to the upper bound on re-identification. Utilizing k-anonymity on eye-tracking feature data, David-John et al. [343] showed that person re-identification accuracies drop to chance levels. In contrast, the utility of a model trained on privacy-preserving data is kept within reasonable ranges when the document-type classification was the utility task, similar to the works that focused on differential privacy. However, the methods used to achieve $k$-anonymity typically depend on the duplication or generalization of data, which is not ideal when releasing datasets for research or statistical purposes. This issue led David-John et al. [344, 343] to introduce the guarantee of $k,\gamma$-plausible deniability to eye-tracking data. Plausible deniability differs from $k$-anonymity in that it applies explicitly to synthetic data generated by a model. All synthetic data generated by the model is tested to ensure that it meets the following privacy guarantee before release.
@@
Plausible Deniability. For any dataset $D$ where $|D|\geq k$, and any record $y$ generated by a probabilistic generative model $\mathbf{M}$ such that $y=\mathbf{M}(d_{1})$ for $d_{i}\in D$, it is stated that $y$ is releasable with ($k$,$\gamma$)-plausible deniability if there exist at least $k-1$ unique records $d_{2},...,d_{k}\in D\setminus\{d_{1}\}$, such that
@@
When considering datasets of gaze samples, the researchers found that different formal methods produced practical privacy-utility trade-offs across different applications. Namely, practical trade-offs were achieved for the utility task of activity type recognition for both plausible deniability and a sample-based differential privacy method, while for a gaze prediction task, only $k$-anonymity produced a dataset that had minimal errors when training a deep neural network. The takeaway from these works is that the recommended method often depends on the target application and whether privacy protections require differential privacy or guarantees specific to re-identification.
In another work, Li et al. [346] adapted a differential privacy approach for providing real-time privacy control of $\epsilon$ for eye-tracking data based on location indistinguishability [347]. The key difference between this approach and the other methods described earlier is the spatial privacy parameter, $r$. The authors proposed tuning the value of $r$ to objects currently in view of the user, which requires an object detection model to be run in parallel to the privacy noise method. The method was applied in real-time to an interactive gaze-controlled action game using webcam-based eye tracking and offline to a VR dataset of video viewing. The authors reported that, according to subjective feedback from their subjects, the subjects enjoyed the gaming experience with real-time privacy protection. Further, while the game scores decrease in high-privacy settings (i.e., $\epsilon=0.5$) compared to no privacy setting, the reduction is modest, with participants still tend to enjoy the game experience. Hence, the privacy-utility, and in this particular case, the privacy-interactivity trade-off, was achieved by finding an empirical balance between the randomly generated noise to protect privacy and asking users whether they enjoy the interaction experience with privacy protection. However, their real-time interaction utility evaluation was limited to 11 participants. In addition, while the suggested method effectively protects against re-identification, the VR dataset application was limited to predicting the users’ visual correction prescription. Table IX summarizes the existing formal privacy methods for eye-tracking data, their corresponding guarantee, data type, and whether the trade-off was considered practical regarding privacy and utility.
TABLE IX: Summary of formal privacy methods applied to eye movement data. The checkmarks indicate whether there is a practical trade-off for the utility task and application.

### v3

@@ IV-D Summary
With eye tracking becoming pervasive, three main directions exist in the literature. The computer vision and machine learning communities address eye region segmentation, gaze estimation, and eye event classification tasks to facilitate accurate attention tracking in VR. Deep learning-based end-to-end approaches are successful when considering various algorithms and methods, particularly for eye region segmentation and gaze estimation. To facilitate the use of eye-tracking technology in an everyday manner in VR, various researchers and device vendors have recently embedded eye trackers in VR HMDs and provided the community with publicly available datasets to facilitate benchmarking and open science.
Once the eye regions and gaze directions are available, gaze-based interaction, including target selection and text entry through eye movements, predictive gaze analyses covering gaze and scanpath prediction, and gaze-contingent rendering, are the most prominent research directions. In most of these works, real-time working capability and eye movement signal quality are essential to achieve a good interaction experience. Therefore, if the gaze estimation and eye-tracker calibration work accurately, it is likely that methods relying on gaze provide a good usability and user experience.
While real-time working capability and eye-tracking data quality are essential for gaze-based interaction in VR, eye movements have also been studied offline to gain a deeper understanding of human cognition, visual attention, and perception. As one can assess these human aspects in various domains in VR, the studies have a broader focus. Prominent domains include education and medicine, as there is a pressing need for online teaching and learning, especially following the COVID-19 pandemic. In these domains, it is possible to control environmental and stimulus factors with precision in VR, which is particularly essential for the medical domain. Regardless of the studied domains, fixation, saccade, blink, pupil diameter, and scanpath statistics are commonly used to assess human cognition, attention, and perception in VR.
@@ V Security and Privacy in Eye Tracking and Implications for Virtual Reality
It is possible to extract and infer useful information about users with eye-tracking data in VR; therefore, eye tracking is considered a powerful modality. We mainly considered works that utilize or are directly applicable to VR. However, virtual environments are relatively novel, and researchers have not evaluated possibilities with eye movements extensively in VR compared to real-world settings. [4] previously stated that, apart from person identification, eye movements can reveal many user attributes, such as gender, sexual preferences, body mass index, health status, or tasks, mainly in real-world settings. Considering that human eye movements depend on the stimulus and context, we first argue that as long as users encounter similar stimuli in virtual environments, comparable inferences will be possible, like in the real world, as some of them, such as gender and task predictions, can already be carried out accurately in VR [18, 279]. We foresee many other inference possibilities in everyday VR.
To this end, [280] stated that while the existence of eye-tracking data is vital for utility tasks such as improving the efficiency of VR rendering, it also gives opportunities, especially for companies, for targeted advertising based on such unique characteristics and demographics of users. Considering these, as identification of users might help personalize VR, we first discuss authentication possibilities in Section V-A. However, if the purpose is not to authenticate users, carrying out the same or similar tasks may result in a privacy breach. The National Institute of Standards and Technology (NIST) defines the privacy breach as “the loss of control, compromise, unauthorized disclosure, unauthorized acquisition, or any similar occurrence where a person other than an authorized user accesses or potentially accesses data or an authorized user accesses data for an other than authorized purpose” [281]. According to the NIST’s definition, access and action by authorized entities or data processing with consent does not constitute a privacy breach. However, with increased digitalization, adversarial entities may cause unintended privacy breaches. Considering this, previous research studied how to preserve data and user privacy. Section V-B covers these topics and discusses methods for preserving privacy when eye tracking is the focus in VR.
@@ V-A Eye-based Authentication for VR
Visual interaction and user assistance are two of the most prominent applications of behavioral data, including eye movements, particularly in real-time VR applications. Apart from those, such behavioral data have also been utilized for authentication purposes. One major challenge for behavioral authentication is that the accuracy of the methods is not very high, which can cause unpleasant user experiences.
@@
Several works in the literature studied eye and gaze-based authentication in VR. [282] studied authentication in VR by utilizing body motion data, including head, hand, and eye movement data for different tasks such as pointing, grabbing, walking, and typing. The results indicate that while the best-achieved identification accuracies are near $60\%$, the performance drops as the user group size increases. However, they did not utilize any gaze features beyond the gaze ray. [283] studied user authentication based on VR spatial movements and reached an accuracy of up to $90\%$. In addition, another study [284] demonstrated that gazing behaviors and head orientation can authenticate users in VR with nearly perfect accuracy. However, the limitation of those studies is the small sample size. Therefore, it remains an open question as to how the utilized features perform when more users are available and how usable the proposed authentication methods would be. That is why more distinguishing and fine-grained features are needed for eye-based authentication. One could achieve this using high-sampling frequency eye trackers.
@@
When we consider eye-based authentication, there are two primary directions: iris-based and eye movement-based authentication. On the one hand, while iris textures are like visual fingerprints and help authenticate users with very high accuracies [285, 286], and iris-based authentication can also be used in real-world scenarios such as border crossings [287]. Due to privacy reasons, many commercial HMDs with eye trackers do not provide raw eye images to users or applications (e.g., HTC Vive Pro Eye [1]). This precaution can somewhat mitigate the problem; however, in the case of iris-textures utilized in VR, data protection issues should be handled carefully, possibly by using encryption [285]. On the other hand, even if the iris textures are not captured or are inaccessible, eye movements will still be provided, making it plausible to carry out authentication using them in the background over a certain period. While we do not conduct an extensive review of eye-based authentication, we provide essential studies related to implicit biometric authentication, particularly utilizing eye movements in VR, and discuss why these studies are relevant to the privacy aspects of such data. For more detailed information on authentication, and particularly eye-based authentication, we refer the reader to the prior survey papers [288, 19].
@@
Task-independent authentication is possible using eye movements [289]. More related to VR, [290] argued that one could use specific eye movement features for biometric authentication in affordable consumer-level devices during everyday tasks such as reading, writing, and web browsing. While the authors did not use a VR setup, their sampling rate of 50 Hz is even lower than the sampling rates of today’s consumer-grade HMDs’ eye trackers. In follow-up work, [291] presented a continuous authentication system based on eye movement biometrics and stated that for eye movement-based authentication, a precise calibration should be carried out and the effects of light sensitivity and task dependence of eye movements should be considered when designing the authentication systems. In another work, [292] proposed a continuous authentication scheme based on eye movements for VR HMDs, demonstrating the potential for HMD personalization by continuously authenticating the wearer in the background. Similarly, [293] used blinks and pupil sizes for user authentication in VR HMDs.
@@
While previous work demonstrates the plausibility of biometric authentication based on eye movements, most of these approaches suffer from usability or privacy issues in the context of VR, which is similar to behavioral authentication using spatial movements. Concerning usability, authentication models based on eye and gaze movements do not have as high accuracies as iris authentication, which could irritate HMD users if errors occur constantly. Furthermore, such models primarily rely on the temporal movement of the gaze, regardless of whether authentication is explicit or implicit, which results in longer authentication times compared to using single iris images. [294] stressed the similar implications and partly confirmed that biometric authentication performance using eye movements is comparable to four-digit PIN entry when eye-tracking signal quality is extremely high, along with high sampling frequencies (i.e., 1kHz). The authors stated that five seconds of eye movement data are needed to achieve such performance. [295] also measured measurement time and presentation size of stimuli on biometric authentication, and they found that using a three-second timeframe maximized the biometric authentication success rates. While the authors did not use a VR setup for this, one can transfer the results to VR as well, as long as the quality of the eye-tracking data and presentation sizes are comparable.
@@
These findings suggest that while implicit eye-based authentication may not offer the same level of usability and security as PIN-based authentication, eye movements exhibit patterns that can help identify individuals to a certain extent. This indication may thus be more useful for personalization, where each user’s data samples are accumulated over time while using VR HMDs and utilized for authentication-related settings such as Two-Factor Authentication (2FA). Since such an authentication process should run continuously in the background, longer processing times would not irritate users, as they would not be actively aware of the ongoing process.
@@
Several works could support this argumentation. For instance, [296] proposed a real-time capable architecture for eye movement-based authentication in VR and argued that eye-based biometrics would become a standard way of authentication for VR. [297] explored the human visual system as a novel authentication method for VR HMDs and proposed OcuLock, which is resistant against impersonation and statistical attacks while maintaining a stable performance over two months, with an increasing error trend in longer timespan. Increased error rates in longer timespan pose another prominent challenge than the authentication accuracy itself, which requires further research. Furthermore, [298] assessed the effectiveness of biometric feature normalization techniques, including real and synthetic eye movement features, and found that the effectiveness of biometric normalization techniques on real-world data depends on feature intercorrelation. [299] used VR to evaluate the usability and security of real-world authentication systems, showing VR’s potential as a testbed for authentication.
@@
Despite several disadvantages, such as the need for consistent tasks, high-sampling frequency eye movement data, and a decrease in authentication performance with a larger sample pool, previous work consistently showed that eye movement patterns represent personal identities, which is useful for authentication and personalization. However, if users do not want to take advantage of these possibilities and yet want to use gaze-based interactions and obtain assistive support, the linkage of eye movements to user identity and characteristics implies privacy risks that should be handled methodologically, meaning that personal patterns associated with eye movements should be either hidden in the data or handled in a privacy-preserving manner. At the same time, one should maintain high interaction utility and user experience.
@@ V-B Privacy-preserving Eye Tracking for VR
Research in eye-based authentication [19], privacy considerations for eye tracking [4, 16], and privacy risks of data collection in XR [300] demonstrate that when authentication is not preferred, it is advisable to preserve the privacy of the individuals and specific attributes of the individuals in the data, in addition to good data hygiene practices. Furthermore, [301] highlighted data privacy issues in VR, considering threat and defense models, and included eye-tracking sensors and data as discussion points. This section covers the methodological research on privacy protection of eye-tracking data. To provide users with privacy-aware eye-tracking solutions, privacy-preserving methods have been applied to sensor data in the form of eye images, aggregated data in the form of saliency maps and feature vectors, and sample-level gaze position data. Each method transforms or modifies data in various ways, creating a trade-off between data privacy and utility that varies by method and the intended application of the data.
Recently, formal methods such as differential privacy [302] have attracted the attention of VR and eye-tracking communities. Differential privacy is an established approach, applied across multiple domains, most prominently in the context of databases and aggregate survey data. It aims to conceal the existence of an individual’s records in a particular database. This privacy mechanism is achieved by adding randomly generated noise, as specified by an $\epsilon$ parameter, to query outputs, ensuring that including a specific participant in the database does not significantly alter the queried function outputs. In $\epsilon$-differential privacy, the smaller the $\epsilon$ is, the more private the mechanism becomes. In the following, we provide the formal definition of $\epsilon$-differential privacy.
@@
$\epsilon$-Differential Privacy ($\epsilon$-DP) [302]. A mechanism $M$ is considered $\epsilon$-differentially private for all databases $D_{1}$ and $D_{2}$ that differ at most in one element for all $S\subseteq Range(M)$ with;
@@
The eye-tracking community initially utilized standard differential privacy mechanisms, namely Gaussian and Exponential mechanisms, on saliency heatmaps [303] and aggregated eye movement features that were collected from VR reading tasks [18]. These works showed the potential of formal methods for VR, even with $\epsilon$ values in the vicinity of $\epsilon=1$; however, they suffer from the correlations in the data that could jeopardize privacy, especially from temporal correlations, as independent noise sampling from the standard mechanisms allows adversaries to reconstruct signals that are very close to the original ones. Despite the correlation issue, both works showed that it is possible to maintain high accuracy in utility tasks while preserving privacy.
@@
Following up on these works, [279] addressed the temporal correlation issue by translating the data into difference signals and using the frequency domain to add the randomly generated noise. The authors noted that, apart from the correlation challenge, the privacy-utility trade-off is particularly important; they also indicated that, despite the challenges, finding an optimal trade-off is computationally feasible within the range $0<\epsilon<1$. Considering that eye trackers may already provide noisy sensor readings due to technical issues, finding a privacy-utility trade-off by adding more noise is an additional challenge. [304] adopted a local differential privacy approach to address the privacy-utility trade-off in a gaze trajectory generation task, arguing that their approach could effectively protect individual user privacy without significant utility loss. Yet, the noise required to establish a strong differential privacy guarantee is known to harm utility and may end up masking valuable insights from the processed data [305]. To address this, the authors of the previous works that studied differential privacy aimed to calibrate the noise in a way that while privacy is achieved in the reasonable ranges (e.g., $\epsilon<1$), they tried to keep the utility tasks (e.g., document type classification) accurate and such tasks performed with accuracies double the amount of chance [279]. Yet, preserving privacy in this manner is not trivial.
@@
Researchers have also explored alternative privacy guarantees beyond differential privacy to gain a deeper understanding of the range of privacy-utility trade-offs associated with eye-tracking data, specifically $k$-anonymity and $k,\gamma$-plausible deniability [306, 307]. These alternative guarantees target the privacy risk of re-identification, and their mathematical formulation is related to the ability to match released data to the original identity in a dataset. Alternative guarantees enable dataset owners to preserve privacy by reducing the risk of linking data to identities, thereby mitigating harms related to the inferences that can be made from eye tracking, while retaining data utility across various applications. First, $k$-anonymity is a definition of privacy in the context of re-identification from a dataset as follows [308].
@@
$k$-anonymity. Given a person-specific dataset $D_{1}$, a de-identified dataset $D_{2}$ is $k$-anonymized by privacy process $\mathcal{P}:D_{1}\mapsto D_{2}$ if all released features $\Gamma_{d}=\mathcal{P}(\Gamma)\in D_{2}$ cannot be recognized as $\Gamma$ with probability $\frac{1}{k}$.
The privacy guarantee is interpreted with a lone privacy parameter $k$ linked to the upper bound on re-identification. Utilizing k-anonymity on eye feature data, [306] showed that person re-identification accuracies drop to chance levels. In contrast, the utility of a model trained on privacy-preserving data remains within reasonable ranges when the document-type classification is the utility task, similar to the works on differential privacy. However, the methods used to achieve $k$-anonymity typically depend on the duplication or generalization of data, which is not ideal when releasing datasets for research or statistical purposes. This issue led [307, 306] to introduce the guarantee of $k,\gamma$-plausible deniability to eye-tracking data. Plausible deniability differs from $k$-anonymity in that it applies explicitly to synthetic data generated by a model. All synthetic data generated by the model is tested to ensure that it meets the following privacy guarantee before release.
@@
[…]
@@
When considering datasets of gaze samples, researchers found that different formal methods yielded practical privacy-utility trade-offs across various applications. Namely, practical trade-offs were achieved for the utility task of activity type recognition using both plausible deniability and a sample-based differential privacy method. In contrast, for a gaze prediction task, only $k$-anonymity produced a dataset with minimal errors when training a deep neural network. The takeaway from these works is that the recommended method often depends on the target application and whether privacy protections require differential privacy or guarantees specific to re-identification.
In another work, [309] adapted a differential privacy approach for providing real-time privacy control of $\epsilon$ for eye-tracking data based on location indistinguishability [310]. The key difference between this approach and the other methods described earlier is the spatial privacy parameter, $r$. The authors proposed tuning the value of $r$ to objects currently in view of the user, which requires an object detection model to be run in parallel to the privacy noise method. The method was applied in real-time to an interactive gaze-controlled action game using webcam-based eye tracking and offline to a VR dataset of video viewing. Subjective feedback from participants indicated that the game was enjoyable with real-time privacy protection. Furthermore, while the game scores decrease in high-privacy settings (i.e., $\epsilon=0.5$) compared to the no-privacy setting, the reduction is modest, with participants still tending to enjoy the game experience. Hence, the privacy-utility and, in this particular case, the privacy-interactivity trade-off were achieved by finding an empirical balance between the randomly generated noise to protect privacy and asking users whether they enjoyed the interaction experience with privacy protection. However, their real-time interaction utility evaluation was limited to 11 participants. In addition, while the suggested method effectively protects against re-identification, the VR dataset application was limited to predicting the users’ visual correction prescription. Table IX summarizes the existing formal privacy methods for eye-tracking data, corresponding guarantee, data type, and whether the privacy-utility trade-off was considered practical.
TABLE IX: Summary of formal privacy methods applied to eye movement data.